# Wigner distribution, Wigner entropy, and Anomalous Transport of a Generalized Aubry-André model

### v2

@@ Abstract
In this paper, we study a generalized Aubry-André model with tunable quasidisordered potentials. The model has an invariable mobility edge that separates the extended states from the localized states. At the mobility edge, the wave function presents critical characteristics, which can be verified by finite-size scaling analysis. Our numerical investigations demonstrate that the extended, critical, and localized states can be effectively distinguished via their phase space representation, specially the Wigner distribution. Based on the Wigner distribution function, we can further obtain the corresponding Wigner entropy and employ the feature that the critical state has the maximum Wigner entropy to locate the invariable mobility edge. Finally, we reveal that there are anomalous transport phenomena between the transition from ballistic transport to the absence of diffusion.
@@ I Introduction
Anderson localization Anderson (1958), a fundamental phenomenon in wave propagation through disordered media, continues to be an active field in condensed matter physics. Scaling theory Abrahams et al. (1979), applied to disordered systems, highlights the critical influence of spatial degrees of freedom on Anderson localization. In one- and two-dimensional systems, the introduction of uncorrelated random disturbances results in the exponential localization of all wave functions. Consequently, the localization-delocalization transition has been a longstanding focus in low-dimensional disordered systems. Conversely, three-dimensional (3D) Anderson system Anderson (1958); Mott (1987) displays a unique behavior, where wave functions are neither entirely localized nor delocalized. An energy threshold, known as mobility edge, separates delocalized states from localized ones. However, detecting the mobility edge in 3D systems experimentally remains difficult. Therefore, to gain a deeper understanding of the mobility edge, low-dimensional systems—particularly one-dimensional (1D) systems—offer a more feasible research path.
Except for the Anderson systems, the 1D Aubry-André (AA) model Aubry and André (1980) with a self duality plays a similarly important role in understanding the Anderson localization and the mobility edges as well. The AA model has an extended-localized transition point, which can be exactly derived by the dual transformation. In addition, AA model is also one of the important source for designing systems with mobility edges. The pursuit of systems with mobility edges has a rich history, including the development and study of numerous theoretical models over the years Abrahams et al. (1979); Mott (1987); Soukoulis and Economou (1982); Lee and Ramakrishnan (1985); Das Sarma et al. (1986, 1988, 1990); Liu et al. (2017a); Biddle and Das Sarma (2010); Biddle et al. (2011); Ganeshan et al. (2015); Liu and Guo (2018); Ye et al. (2024). Furthermore, recent experiments have successfully identified mobility edges in one-dimensional quasi-periodic optical lattices by manipulating cold atoms Lüschen et al. (2018); An et al. (2018); Kohlert et al. (2019). Specifically, Das Sarma et al. revealed that there was mobility edge in a class of systems with slow-varying quasidisordered on-site potentials Das Sarma et al. (1988, 1990); Liu et al. (2017a). Biddle et al. found the exactly solvable mobility edge in a generalized AA model with long-range tunnelings Biddle and Das Sarma (2010); Biddle et al. (2011). Ganeshan et al. verified the existence of the mobility edge protected by the self duality Ganeshan et al. (2015). Liu et al. obtained the exactly solvable mobility edge in the off-diagonal quasidisordered models Liu and Guo (2018). Wang et al. uncovered a class of exactly solvable 1D models with mobility edgesWang et al. (2020). Zhou et al. not only proposed a class of exactly solvable models which host a novel type of exact mobility edge, but also propose a novel experimental scheme to realize the exactly solvable model and the new mobility edges in an incommensurate Rydberg Raman superarray Zhou et al. (2023). Wang et al. obtained the analytical expression of the anomalous mobility edge, which separates localized states from critical states in the case with unbounded quasiperiodic potential Wang et al. (2023). Guan et al. studied localization properties and mobility edges of a generalized spinful Aubry-André-Harper (AAH) model, and analytically obtained exact expressions of the mobility edges by a semiclassical method.Guan et al. (2023). Liu investigated the properties of mobility edges in 1D incommensurate lattices with p-wave superfluidity Liu et al. (2022a). Ye et al. studied the mobility edge in a family of quasiperiodic systems evolving far from equilibrium Ye et al. (2024). Yamilov et al. investigated the Anderson transition for 3D light and identified a mobility edge that separates diffusive transport and Anderson localization Yamilov et al. (2025). Longhi demonstrated the existence of a mobility edge in quasicrystals induced by pure dephasing effects through the study of photonic quantum walks in synthetic mesh lattices Longhi (2024), and also provided an exact spectral analysis of the mosaic Wannier-Stark Hamiltonian, showing that there are no mobility edges strictly separating extended and localized states Longhi (2023). Qi et al. discovered quantum Griffiths singularities in ${\rm MgTi_{2}O_{4}}$, where the mobility edge of Anderson localization may lead to the occurrence of a 3D quantum phase transition and the emergence of quantum Griffiths singularities Qi et al. (2024). Chang et al. experimentally investigated a mosaic quasiperiodic photonic waveguide lattice and discovered the mobility edge Chang et al. (2025). Gonçalves et al. proposed a class of solvable one-dimensional quasi periodic tight binding models, in which the extended phase, localized phase, and critical phase are separated by mobility edges Gonçalves et al. (2023).
In addition to analytical demonstrations and experimental observations, the extended-localized transition can be observed by the dynamical evolution as well. Xu et al. have studied the wave packet dynamics of a one-dimensional system with generalized quasidisordered potential and parameter-dependent mobility edge Xu et al. (2020), and found that in the extended regime and the regime with mobility edge, the wave packet spreads with the behavior of ballistic transport, while in the localized regime, the wave packet does not diffuse. In fact, there is anomalous spreading phenomenon. It means that, as the change of systematic parameter, the spreading form of the wave packet may successively exhibit phenomena such as ballistic transport, superdiffusion, diffusion, subdiffusion, and non-diffusion Purkayastha et al. (2021). For example, for the Fibonacci model whose on-site potential is arranged with a Fibonacci substitution rule Chiaracane et al. (2021) and the AA model in isolated and open systems Purkayastha et al. (2018); Lacerda et al. (2021), there are anomalous transport phenomena. These findings drive us to inquire whether similar anomalous transport phenomena can occur in a system with a generalized quasidisordered potential and invariable extended-localized mobility edge. In 1932, Eugene Wigner preposed the phase space representation, i.e., Wigner distribution function $W(x,p)$ (here $x$ and $p$ denote the coordinate and momentum, respectively), which enables the concurrent description of quantum states or signals in both coordinate and momentum Wigner (1932); Wig (1984); Kenfack and Życzkowski (2004); Wig (2014); Simon (2000); Taghiabadi et al. (2016); Siyouri et al. (2016). It serves as a distinctive “fingerprint” that captures the joint distribution of these variables, hence providing a unique representation for quantum states or signals. This prompts us to further reveal the difference between the extended state and the localized state in both coordinate and momentum dimensions, and to further consider how to employ the Wigner distribution function to locate the mobility edge.
@@ II Model and the mobility edge
where $\hat{c}^{{\dagger}}_{n}$ ($\hat{c}_{n}$) is the creation (annihilation) operator with $n$ the site index, $L$ is the length of the system and $t$ is the hopping strength which is set as the unit of energy. $V_{n}=\frac{V}{1-b\cos(2\pi\alpha n)}$ is the generally quasidisordered potential, in which $V$ is the strength of the potential with $10<b<$ 1 being the tuning parameter and $\alpha=\frac{\sqrt{5}-1}{2}$ is the incommensurate parameter. In coordinate space, we can define the generalized wave function |$\ket{\psi}=\sum_{n}\phi_{n}\hat{c}^{{\dagger}}_{n}\ket{0}$ with $\phi_{n}$ the amplitude of probability. Based on |$\ket{\psi}$, we can obtain the following static Schrödinger equation
@@
Employing the self-dual transformation, we can obtain the exact expression of the extended-localized mobility edge $E_{c}$ of the generalized AA model Liu et al. (2022b), which satisfying $E_{c}=\frac{2t}{b}$. From the exact expression of $E_{c}$, we know that the mobility edge is invariant with $V$, but is uniquely determined by $b$. In the following, we will show that in addition to the inverse participation ratio and finite-size scaling analysis to study the extended-localized transition over the energy domain, the extended-localized transition can be characterized by the Wigner distribution as well. Based on the Wigner distribution function, we can further define the winger entropy and show that the mobility edge can be located by winger entropy, and the results are consistent with the analytical solution. Finally, we will show that the system with generalized quasidisordered potential has the property of anomalous quantum transport.
@@
Taking $b=0.4$ and $L=500$, the resulting localization phase diagram is plotted in Fig. 1. The color bar denotes the value of inverse participation ratio (IPR) defined as ${\rm IPR}_{j}=\sum^{L}_{n=1}|\phi^{j}_{n}|^{4}$ with $j$ the energy level index. For an extended state, ${\rm IPR}\propto L^{-1}$ and decays to zero under larger $L$. For a localized state, ${\rm IPR}$ is a finite value Biddle and Das Sarma (2010); Biddle et al. (2011). From Fig. 1, we can see that the invariable mobility edge $E_{c}/t=5$ perfectly separates the extended states from the localized state. Below $E_{c}$ ${\rm IPR}$ tends to 00, signaling the extended states. Above $E_{c}$, ${\rm IPR}$ are finite values, signaling the localized states. The numerical results are consistent with analytical results. To be specific, we plot the typically extended and localized state in Figs. 2(a) and 2(d), whose energy is far below $E_{c}$ and far above $E_{c}$, respectively. We can see that the probability distribution of the extended state extends over all the system while that of the localized state only occupies a small part of the system. In addition, at $E_{c}$, the wave functions are critical, because we can see from Figs. 2(b) and 2(c) that although the probability distribution occupies most of system space, there is still a part of the system space that is not occupied, presenting the characteristics of subexpansion.
[…]
Having recalled the extended-critical transition and the invariable caused by the generalized potentials, in the following section, we will focus on employing the Wigner distribution, a phase space pattern, to distinguish extended states and the localized states, and employing Wigner entropy to locate the mobility edge. In addition, we will reveal the anomalous transport phenomenon of the model.
@@ III Wigner distribution and Wigner entropy
For the given wave function |$\ket{\psi}$, the Wigner distribution function $W(x,p)$ can be obtained by the following integration Wigner (1932); Wig (1984); Kenfack and Życzkowski (2004); Wig (2014); Simon (2000); Taghiabadi et al. (2016); Siyouri et al. (2016)
@@
Based on the Wigner distribution, one can further obtain the Wigner entropy (marked by $W_{S}$) Van Herstraeten and Cerf (2021). Considering the negativity of $W(x,p)$, the $W_{S}$ is finally calculated by the following definition
@@
To verify our inference, we plotted $W_{S}$ for $V=3.2t$, $V=4.8t$, and $V=7.3t$ in Figs. 4(a), 4(b), and 4(c), respectively. From the figures, we can see that $W_{S}$ peaks at $E_{c}=5t$ which means that the critical state has a maximal $W_{S}$ at $E_{c}$. Moreover, the $W_{S}$ below $E_{c}$ is slightly larger than that above $E_{c}$. Concretely, the maximal $W_{S}$ for the $V=4.8t$ case is $W^{critical}_{S}=5.2217$. Besides, we make averages on $W_{S}$ before and after $E_{c}$, respectively, and obtain the averaged $W_{S}$ for the extended states satisfies $W^{extended}_{S}=3.7875$ and that for the localized states is $W^{localized}_{S}=2.6249$. From the numerical results, it is intuitive that $W^{critical}_{S}>W^{extended}_{S}>W^{localized}_{S}$. Therefore, the inference made from the distributions of $W(x,p)$ is correct under the validations of numerical calculation. We emphasize that this conclusion is universal to the results for other $V$ and $b$. In addition, the results tell that we can locate the mobility edge by using the property that the mobility edge produces the maximum Wigner entropy, and distinguish the extended state from the local state by comparing the average entropy of the left and right sides of the maximum. Meanwhile, a full phase diagram that contains $W_{S}$ in the $E$-$V$ parameter space has been presented in Fig. 4(d). The color bar represents $W_{S}$, and from the color point of view, the mobility edge can separate the two states of extension and localization. Recently, there are different ways to realize the measurement the Wigner distribution function, such as the qubit quantum processor Tilma et al. (2016); Rundle et al. (2017), discrete atomic system Tian et al. (2018), photonic system Tilma et al. (2022); Xu et al. (2014), and optical parametric amplification Kalash and Chekhova (2023). Wigner entropy is a statistical result of Wigner distribution. With these experimental techniques, it may be possible to experimentally locate the mobility edge accurately in the future.
Figure 5: (Color Online) The time evolution of mean squared displacement $\Delta x^{2}(\tau)$ under $b=0.4$ and $L=1000$. (a) $V=0.2t$; (b) $V=2.5t$; (c) $V=4.8t$; (d) $V=6.8t$; (e) $V=7.3t$; (f) $V=10t$.

### v3

@@ Abstract
We investigate generalized Aubry-André models featuring tunable quasidisordered potentials and a mobility edge that separates extended and localized states, with critical states for the mobility edge confirmed through finite-size scaling analysis. Numerical results demonstrate that extended, critical, and localized states can be distinguished via their phase-space representations, particularly the Wigner distribution. The associated Wigner entropy, derived from this distribution, peaks at the critical state, enabling precise localization of the mobility edge. Additionally, wave-packet dynamics reveal anomalous transport behaviors, including superdiffusion and subdiffusion, bridging ballistic transport and the absence of diffusion.
@@ I Introduction
Anderson localization [1], a fundamental phenomenon in wave propagation through disordered media, continues to be an active field in condensed matter physics. Scaling theory [2], applied to disordered systems, highlights the critical influence of spatial degrees of freedom on Anderson localization. In one- and two-dimensional systems, the introduction of uncorrelated random disturbances results in the exponential localization of all wave functions. Consequently, the localization-delocalization transition has been a longstanding focus in low-dimensional disordered systems. Conversely, three-dimensional (3D) Anderson system [1, 3, 4] displays a unique behavior, where wave functions are neither entirely localized nor delocalized. An energy threshold, known as mobility edge, separates delocalized states from localized ones. However, detecting the mobility edge in 3D systems experimentally remains difficult. Therefore, to gain a deeper understanding of the mobility edge, low-dimensional systems—particularly one-dimensional (1D) systems—offer a more feasible research path.
Except for the Anderson systems, the 1D Aubry-André (AA) model [5] with a self duality plays a similarly important role in understanding the Anderson localization and the mobility edges as well. The AA model has an extended-localized transition point, which can be exactly derived by the dual transformation. In addition, AA model is also one of the important source for designing systems with mobility edges. Biddle and Das Sarma introduced long-range transitions in the AA model [6, 7], deriving exact expressions for the mobility edge via a self-dual transformation and verifying them numerically through calculations of the inverse participation ratio. This approach established a paradigm for analytically determining mobility edges in subsequent studies using self-duality [8, 9, 10, 11, 12], with numerical validation often employing the inverse participation ratio in generalized AA models [13, 14, 15, 16, 17]. Mobility edges also appear in slowly varying systems, where they can be obtained through semiclassical methods [18, 19, 14, 9]. More recently, mobility edges have been identified in AA systems with mosaic modulation [20], with exact solutions derived using Avila’s global theory [21, 22]. This analytical framework has since been extended to generalized mosaic models [23, 24] and spinful systems [25]. Advances in experimental techniques have enabled the observation of mobility edges via wave-packet dynamics measurements [26, 27, 28, 29].
In the context of wave-packet dynamics, the spreading behavior of wave packets can exhibit five distinct transport regimes: ballistic transport, superdiffusion, diffusion, subdiffusion, and localization [30, 31, 32, 33]. For a generalized AA model featuring a mobility edge separating extended and localized states [16], prior work has demonstrated that wave packets propagate ballistically in both the extended regime and near the mobility edge, whereas they remain confined without diffusion in the localized regime. In addition to ballistic transport, diffusion, and localization, two anomalous transport behaviors—superdiffusion and subdiffusion—have also been identified [30]. These findings motivate an exploration of whether such anomalous transport phenomena can arise in systems with generalized quasi-disordered potentials and extended-localized mobility edges. In 1932, Eugene Wigner preposed the phase space representation, i.e., Wigner distribution function $W(x,p)$ (here $x$ and $p$ denote the coordinate and momentum, respectively), which enables the concurrent description of quantum states or signals in both coordinate and momentum [34, 35, 36, 37, 38, 39, 40]. It serves as a distinctive “fingerprint” that captures the joint distribution of these variables, hence providing a unique representation for quantum states or signals. The Wigner distribution function simultaneously encodes information about a quantum state in both position and momentum spaces, a capability not afforded by the inverse participation ratio. Many mathematical methods can be used to convert the experimental data from optical homodyne tomography into the state’s Wigner function [41] , which makes it possible to experimentally measure the Wigner function. This prompts us to further reveal the difference between the extended state and the localized state in both coordinate and momentum dimensions, and to further study how to employ the Wigner distribution function to locate the mobility edge.
@@ II Model and the mobility edge
[…]
@@
Employing the self-dual transformation, we can obtain the exact expression of the extended-localized mobility edge $E_{c}$ of the generalized AA model [42], which satisfying $E_{c}=\frac{2t}{b}$. From the exact expression of $E_{c}$, we know that the mobility edge is invariant with $V$, but is uniquely determined by $b$. In the following, we will show that in addition to the inverse participation ratio and finite-size scaling analysis to study the extended-localized transition over the energy domain, the extended-localized transition can be characterized by the Wigner distribution as well. Based on the Wigner distribution function, we can further define the winger entropy and show that the mobility edge can be located by winger entropy, and the results are consistent with the analytical solution. Finally, we will show that the system with generalized quasidisordered potential has the property of anomalous quantum transport.
@@
Taking $b=0.4$ and $L=500$, the resulting localization phase diagram is plotted in Fig. 1. The moderate $b$ can present the extended local transformation and mobility edge more intuitively. The color bar denotes the value of inverse participation ratio (IPR) defined as ${\rm IPR}_{j}=\sum^{L}_{n=1}|\phi^{j}_{n}|^{4}$ with $j$ the energy level index. For an extended state, ${\rm IPR}\propto L^{-1}$ and decays to zero under larger $L$. For a localized state, $\rm IPR$} is a finite value [6, 7]. From Fig. 1, we can see that the invariable mobility edge $E_{c}/t=5$ perfectly separates the extended states from the localized state. Below $E_{c}$ $\rm IPR$} tends to 0, signaling the extended states. Above $E_{c}$, $\rm IPR$} are finite values, signaling the localized states. The numerical results are consistent with analytical results. To be specific, we plot the typically extended and localized state in Figs. 2(a) and 2(d), whose energy is far below $E_{c}$ and far above $E_{c}$, respectively. We can see that the probability distribution of the extended state extends over all the system while that of the localized state only occupies a small part of the system. In addition, at $E_{c}$, the wave functions are critical, because we can see from Figs. 2(b) and 2(c) that although the probability distribution occupies most of system space, there is still a part of the system space that is not occupied, presenting the characteristics of subexpansion.
We can further verify the preliminary judgment of the spatial distribution of wave function properties at fixed system size by finite-size scaling analysis. Here, the finite-size scaling analysis is carried out by calculating the fractal dimension $\beta_{\rm min}$. The properties of the wave function are judged by the results of the fractal dimension under the extrapolation limit of the large system size. The $\beta_{\rm min}$ can be calculated by the fractal theory [43, 44, 45, 46, 47, 48]. As first, we choose a system whose size $L$ is equal to the $m$th Fibonacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by the ratio of two neighboring Fibonacci numbers, i.e., $\alpha=\frac{F_{m}}{F_{m+1}}$. Then we can extract a scaling index $\beta^{m}_{n}$ from the probability $P^{m}_{n}=|\phi^{m}_{n}|^{2}$ by $P^{m}_{n}\sim(1/F_{m})^{\beta^{m}_{n}}$. According to the fractal theory, we know that for an extended state, the maximum of $P^{m}_{n}$ follows the scaling ${\rm max}(P^{m}_{n})\sim(1/F_{m})^{1}$, i.e., $\beta_{min}=1$, and for a localized state, ${\rm max}(P^{m}_{n})\sim(1/F_{m})^{0}$, signaling $\beta_{\rm min}=0$, while for the critical state, whose $\beta_{\rm min}$ is within the interval (0, 1). Taking $V=4.8t$ and a Fibonacci sequence (system sizes), we plot the $\beta_{\rm min}$ of the selected extended (chosen from the lowest eigenstate), critical (chosen at $E_{c}$), and localized states (chosen form the highest eigenstate) as a function of $1/m$ in Fig. 2(e). We can see that under extrapolation limit ($m\rightarrow\infty$), $\beta_{\rm min}$ equals to $1$ for the extended state, and is a finite number within $(0,\leavevmode\nobreak\ 1)$ for the critical state, and equals to 0 for the localized state. Therefore, the finite-size analysis validates the preliminary judgement obtained from the spatial distribution of the wave functions.
Having recalled the extended-critical transition and the invariable mobility edge caused by the generalized potentials, in the following section, we will focus on employing the Wigner distribution, a phase space pattern, to distinguish extended states and the localized states, and employing Wigner entropy to locate the mobility edge. In addition, we will reveal the anomalous transport phenomenon of the model.
@@ III Wigner distribution and Wigner entropy
[…]
@@
Based on the Wigner distribution, one can further obtain the Wigner entropy (marked by $W_{S}$) [49]. Considering the negativity of $W(x,p)$, the $W_{S}$ is finally calculated by the following definition
@@
To verify our inference, we plotted $W_{S}$ for $V=3.2t$, $V=4.8t$, and $V=7.3t$ in Figs. 4(a), 4(b), and 4(c), respectively. From the figures, we can see that $W_{S}$ peaks at $E_{c}=5t$ which means that the critical state has a maximal $W_{S}$ at $E_{c}$. Moreover, the $W_{S}$ below $E_{c}$ is slightly larger than that above $E_{c}$. Concretely, the maximal $W_{S}$ for the $V=4.8t$ case is $W^{critical}_{S}=5.2217$. Besides, we make averages on $W_{S}$ before and after $E_{c}$, respectively, and obtain the averaged $W_{S}$ for the extended states satisfies $W^{extended}_{S}=3.7875$ and that for the localized states is $W^{localized}_{S}=2.6249$. From the numerical results, it is intuitive that $W^{critical}_{S}>W^{extended}_{S}>W^{localized}_{S}$. Therefore, the inference made from the distributions of $W(x,p)$ is correct under the validations of numerical calculation. We emphasize that this conclusion is universal to the results for other $V$ and $b$. In addition, the results tell that we can locate the mobility edge by using the property that the mobility edge produces the maximum Wigner entropy, and distinguish the extended state from the local state by comparing the average entropy of the left and right sides of the maximum. Meanwhile, a full phase diagram that contains $W_{S}$ in the $E$-$V$ parameter space has been presented in Fig. 4(d). The color bar represents $W_{S}$, and from the color point of view, the mobility edge can separate the two states of extension and localization.
However, the use of Wigner entropy to locate mobility edges is not an isolated case; for example, it can also be applied to the Biddle-Das Sarma model [6]. This model represents a generalized AA system with long-range hopping, described by the Hamiltonian
$\hat{H}=\sum_{n,n^{\prime}\neq n}te^{-p|n-n^{\prime}|}\left(\hat{c}^{{\dagger}}_{n+n^{\prime}}\hat{c}_{n}+\hat{c}^{{\dagger}}_{n}\hat{c}_{n+n^{\prime}}\right)+V_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n},$ (5)
where $p>0$, the on-site potential is $V_{n}=V\cos(2\pi\alpha n)$, with $\alpha$ an irrational number, and $t$ is the hopping strength (set as the energy unit).
Figure 5: (Color Online) (a) Energy spectrum $E$ as a function of $V/t$, with the color bar indicating the value of the IPR; (b)-(d) Wigner entropy $W_{S}$ as a function of eigenenergy $E$ for (b) $V=0.266t$, (c) $V=0.413t$, and (d) $V=0.811t$. (e) $W_{S}$ in the $E$-$V$ parameter space, with the solid black line denoting the mobility edge $E_{c}=V\cosh(p)-t$. Other parameters are $p=1.5$ and $L=500$.
Figure 5 presents the energy spectrum $E$ of the Biddle-Sarma model as a function of $V/t$ for $p=1.5$ and $L=500$, with the color bar indicating the IPR. The blue solid line marks the invariant mobility edge $E_{c}=V\cosh(p)-t$ [6]. The Wigner entropies $W_{S}$ are depicted for $V=0.266t$, $V=0.413t$, and $V=0.811t$ in Figs.  5(b), 5(c), and 5(d), respectively. In each instance, $W_{S}$ reaches a peak at the mobility edge $E_{c}$, indicating that the critical state exhibits a maximum $W_{S}$ value there. Moreover, the plots demonstrate that the average Wigner entropy adheres to $W^{critical}_{S}>W^{extended}_{S}>W^{localized}_{S}$. These numerical observations affirm that the Wigner distribution and entropy can effectively discriminate among extended, critical, and localized states, and that the entropy maximum serves as a robust indicator for pinpointing the mobility edge. Recently, there are different ways to realize the measurement the Wigner distribution function, such as the qubit quantum processor [50, 51], discrete atomic system [52], photonic system [53, 54], and optical parametric amplification [55]. Wigner entropy is a statistical result of Wigner distribution. With these experimental techniques, it may be possible to experimentally confirm our theoretical predictions in the near future.